# Bridging Online Behavior and Clinical Insight: A Longitudinal LLM-based Study of Suicidality on YouTube Reveals Novel Digital Markers

### v1

@@ Abstract
Suicide remains a leading cause of death in Western countries, underscoring the need for new research approaches. As social media becomes central to daily life, digital footprints offer valuable insight into suicidal behavior.
Focusing on individuals who attempted suicide while uploading videos to their channels, we investigate: How do suicidal behaviors manifest on YouTube, and how do they differ from expert knowledge? We applied complementary approaches: computational bottom-up, hybrid, and expert-driven top-down, on a novel longitudinal dataset of 181 YouTube channels from individuals with life-threatening attempts, alongside 134 control channels. [^2] Code and data will be available on demand, due to data sensitivity.
[…]
@@ 1.   Introduction
Suicide remains one of the most significant public health challenges, particularly among young people in western countries, where it consistently ranks as one of the leading causes of death [34, 17]. Traditionally, research on suicidality has focused on factors such as mental health issues (e.g., depression), social isolation, and family dynamics [13, 39, 14]. These factors were predominantly predefined by experts and were typically examined by relying on patients’ self-reports. However, as social media becomes increasingly prevalent and suicide rates continue to grow, established risk factors and their manifestation are changing, requiring further investigation into suicidal behavior in digital spaces [9, 40].
@@
Social media have become integral to communication and social development, especially among adolescents [35]. Many young people turn to these platforms to share their experiences and seek support [16, 24], potentially providing a digital footprint, documenting the nuances of everyday life, including moments of joy and distress.
@@
The intersection of suicide, social media, and Artificial Intelligence (AI) has gained traction in recent years, with much research utilizing natural language processing (NLP) to analyze large volumes of social media data [37, 25]. However, a common limitation is a reliance on unvalidated suicide risk assessments, such as identifying risk based on messages like "I want to kill myself" or user participation in suicide-related sub-reddits, which introduces biases in both data collection and reliability. Additionally, many studies lack model transparency and have inadequate control groups. Perhaps most importantly, they often fail to incorporate longitudinal analyses that are essential for capturing temporal shifts in relation to suicidal behavior.
@@
Although social media platforms (e.g., Facebook or Reddit) are commonly used to study suicidality, video-sharing platforms like YouTube remain underexplored. As one of the most widely used multimedia-sharing platforms, YouTube enables users to communicate visually and verbally, providing a more direct and unmediated expression. Moreover, YouTube plays a significant role in both shaping and reflecting societal trends related to mental health and suicidality [2, 21]. Still, a search on PubMed revealed only 28 articles containing the words "suicide" and "YouTube" in their title or abstract, with most focusing on the impact on viewers rather than on the creators [11, 22].
@@
Figure 1: An illustration of our research framework, showing the methodological approaches used in this study. Each circle represents a different approach, with the specific methods listed inside. The computational approach includes LLM-based topic modeling to identify behavioral patterns on YouTube (i. §3.1), analyzed through a psychological lens in the discussion section (§4). The domain-expert approach involves a clinical assessment of suicide attempt narratives of predefined psychological factors (ii. §3.3). In the hybrid approach, a domain expert reviewed all the LLM-derived topics and flagged those that may be considered as suicide risk factors, based on current clinical knowledge (iii. §3.4).
@@
This study addresses existing research gaps by bridging the understanding of suicidality in offline and online contexts, considering both behavioral indicators and self-reported experiences, and by incorporating longitudinal analysis of the development of behaviors and feelings before and after suicide attempts. By applying three complementary approaches, as described in Figure 1 and leveraging the strengths of both computational and domain-expert frameworks (Figure 2), we aim to explore our primary research question: How do suicidal behaviors manifest on YouTube, and how are they similar or different from the clinical literature?
Figure 2: Conceptual comparison between the computational approach and the domain-expert approach as applied in this article.
@@
To allow this investigation, we first constructed a novel dataset of YouTube channels belonging to individuals who had attempted suicide (Figure 3), with cases verified as life-threatening by a clinical psychologist (Figure 4). We focused on individuals who attempted suicide after establishing their YouTube channel, referred as Attempted (During). Alongside the Attempted (During) group, which we defined as the treatment group, we gathered three control groups representing varying levels of suicide risk.
@@
Figure 3: The research groups are categorized into treatment - Attempted (During) - and three control groups - Attempted (Before), Control (Matches) and Control (Major Life Event). Both Attempted groups refer to those who attempted suicide during or before their channel upload period, while the other groups refer to control-based matches from the general population, or to those who experienced a major life event during their upload period (and didn’t attempt suicide as far as we know). The sample size for each group is indicated below their respective categories.
@@
In the bottom-up approach, we employed LLM-based topic modeling and analyzed the topic distributions of each video to identify behavioral patterns on YouTube. By examining differences in pre-suicide attempt average topic probability in the treatment group compared to the control groups, we identified potential indicators of suicidal behavior. Then we performed a longitudinal analysis on the identified topics, revealing temporal trends before and after suicide attempts. Through this analysis, we answer the first part of the research question - how suicidal behaviors manifest on YouTube, and partially address the second part, how are they similar or different from the clinical literature, by examining our findings in light of previous suicide research (see Discussion §4).
@@
Our bottom-up, fully automated computational analysis identified five topics as potential pre-attempt indicators, two of which also exhibited significant temporal shifts around the attempt (§3.1): "Mental Health Struggles" and "YouTube Engagement" (see topics’ representations: Figure 7). "Mental Health Struggles" includes discussions about suicide, self-harm, and depression, aligned with established risk factors. “YouTube Engagement”, which captures typical platform behavior such as audience interaction and content promotion, may serve as an indicator of “normal” YouTube activity, reflecting routine channel maintenance [49, 10, 3]. Notably, this topic appeared at lower levels in the treatment group prior to the suicide attempt—possibly signaling abnormal behavior and, therefore, a potential risk indicator. Both topics exhibited significantly higher prevalence following the suicide attempt, which we discuss at §4.
Additional three topics emerged as pre-attempt indicators and were not associated with a temporal pattern: (1) "LGBTQ+ Identity and Acceptance", (2) "Rape and Sexual Assault", and (3) "Fashion and Style", detailed in §C.3 and discussed in §4. The first two align with established suicide literature, reflecting the heightened distress faced by LGBTQ+ individuals due to discrimination and rejection [46], and by survivors of sexual violence due to trauma-related mental health challenges [12]. Both topics were more prevalent in the attempted groups compared to the other controls. In contrast, Fashion and Style, a less familiar factor in suicide research, showed significantly lower prevalence in the attempted groups, potentially reflecting diminished engagement with self-expression or societal norms.
Moreover, the top-down domain-expert approach showed strong alignment in risk and resilience factors between YouTubers who attempted suicide before and during their upload period, reducing suicide-related tendencies confounders in the comparison between these groups (e.g., depression, family support). The only significant difference was the reason for sharing their attempt story: those who attempted suicide before emphasized their Desire to Help Others, while those who attempted during framed it as part of their Personal Recovery Process (Table 4).
@@ 2.   Methods
This section outlines the methods used in this study, beginning with data collection (Figure 4) and followed by a description of the procedures (Figure 6). The data collection2.1.1) involved constructing a textual dataset comprising four types of YouTube channels creators (Figure 3): belonging to creators who attempted suicide (before or during their YouTube activity), those who experienced a major life event, and matched control users from the general population. We searched YouTube for relevant channels, transcribed the videos and collected channel information, annotated demographic and psychological assessments, and matched them to control channels using a multi-step approach. The procedures2.2) included applying an LLM-based topic modeling to the channels’ transcripts and implementing statistical procedures to address our research objective: understanding how suicidal behavior manifests on YouTube.
@@ 2.1.1 Data collection
YouTube served as our primary data source. We identified users who uploaded a video sharing either their own suicide attempt story or a major life event using the YouTube API. [^4] API Documentation: developers.google.com/youtube/v3/docs These stories were typically not the main theme of the channels, but were often shared to raise awareness of suicide or loss. Additionally, we included a control group that matched the treatment users to facilitate a fair comparison with the general population. The data collection procedure is outlined in Figure 4 and described in more detail below, with full documentation—including inclusion and exclusion criteria—provided in the Supplementary Materials (Table A.1).
Figure 4: The data collection pipeline includes searching for relevant YouTubers (e.g., YouTubers who attempted suicide), a two-step human evaluation, involving non-expert assessment (personal channel validation and extracting demographic and event-related information), and psychological assessment (attempt intent validation and suicide-related factor extraction); both steps validating and then analyzing the narrative stories describing a suicide attempt or a major life event. The final steps involved control-based matching algorithm, which involved finding "similar" YouTube channels to the treatment group (Figure 5). The last step included performing data processing (including filtering uploads and automatic transcription) on the 4 research groups.
@@
We then conducted a human evaluation involving two levels of expertise. First, non-expert channel validation to ensure that only personal channels were included, while excluding educational, news, and family channels. Following this, we extracted user and event information from the attempt and major life event stories (e.g., attempt date). Second, a psychological assessment of the attempt narrative videos of YouTubers who reported a suicide attempt for two purposes: (1) validating the intent of the attempt, and (2) comparing Attempted (During) and Attempted (Before) to suicide-related tendencies to estimate potential confounding factors (e.g., mental illness, risk groups) and to perform a suicide-related factor analysis.
Next, we applied a multi-step matching process (Figure 5), which allowed us to first search for control cases from the general YouTube population and later match them to users who attempted suicide, based on demographic and channel characteristics.
Finally, we employed data processing, which included collecting the users’ channel videos and metadata (e.g., likes, description) and performing automatic transcription using Whisper [41], the current speech-to-text state-of-the-art model.
Figure 5: The matching algorithm, comprising of four steps: (1) creating queries based on the treatment group, using the YouTube API (§A.4.1); (2) automatic and manual filtering resulting in relevant personal channelsA.4.2); (3) propensity score-based matching based on demographic and channel data, using (a) logistic regression and (b) KNN matching of treatment to control users (§A.4.3); and (4) LLM-based refined matching, using GPT-4 to score the potential matching, retaining good matches (scores > 3) (§A.4.4)
@@ 2.1.2 Study population
We categorized channels into two primary groups: treatment and control (Figure 3). The treatment consisted of YouTubers who attempted suicide during their upload period, while the first control group included YouTubers who attempted suicide before their upload period. These groups were labeled as Attempted (During) and Attempted (Before). The additional control groups comprised YouTubers who experienced a major life event during their upload period, labeled Control (Major Life Event), and YouTubers matched to the treatment group based on demographic and content characteristics, labeled Control (Matches). The control groups consist of both a potential at-risk population that had encountered major life events or attempted suicide in the past, and a general population resembling the treatment users. See Table A.1A.1) for inclusion and exclusion criteria.
Group # Users Uploads (Total) Uploads
(Mean / Med / SD)
Female
(n, %)
Age (Mean) Minority (n, %)
Attempted (During) 42 6,354 151 / 89 / 169 30 (71%) 24 27 (59%) Attempted (Before) 139 28,053 201 / 83 / 312 91 (65%) 26 78 (56%) Control (Major Life Event) 43 13,470 320 / 155 / 372 29 (69%) 27 15 (35%) Control (Matches) 95 14,860 182 / 95 / 223 79 (81%) 23 52 (54%)
Table 1: Descriptive Statistics of the Research Groups. Summary of demographics and YouTube activity. See §A.6 for upload trends pre- and post-event.
The Final Sample (Table 1): Of the initial 303 YouTubers in the Attempted group, 104 users were excluded due to non-suicidal cases (e.g., clickbait-like videos like "Bust Myths About Suicide"), non-English content, or an insufficient number of videos, and 18 users were reclassified as Control (Major Life Event) due to the suicide attempt of a close relative. The final set of attempted channels consists of 181 channels (66% female, average age = 25) who uploaded 34,407 videos ($M=190$, $Mdn=83$, $SD=286$). This dataset was later divided into two groups: 42 Attempted (During), defined as the treatment, and 139 Attempted (Before), defined as one of the three controls, based on whether the attempt occurred during or before their upload period (Figure 3).
@@
The final Control (Matches) group included 95 channels (81% female, average age = 23) that met the inclusion criteria and were found as most similar to the 42 Attempted (During) group according to the matching algorithm (Figure 5). These 95 users collectively uploaded 14,860 videos ($M=182$, $Mdn=95$, $SD=223$).
@@ 2.2 Procedures
Figure 6: The methodological pipeline consists of three approaches: 1. Computational Bottom-Up - Factors were generated using an LLM-based topic modeling algorithm, and topic relevance was evaluated using a bottom-up approach; 2. Domain-Expert Top-Down - Factors were predefined by a clinical expert and assessed solely based on the suicide attempt narratives, and a 3. Hybrid Approach - A set of 166 LLM-generated topics was reduced to 19 based on their relevance to suicide as determined by a clinical expert. The bottom-up and hybrid approaches were also followed by a two-way mixed effect ANOVA examining both temporal and group effects, while the top-down approach was limited to group differences.
LLM-based Topic A topic is defined as a cluster of semantically related text segments gathered by an unsupervised clustering method of sentence embeddings. Each topic is automatically titled by LLM considering top-associated words and representative texts.
For topic modeling details, see §B.1.
First, we applied an LLM-based topic modeling technique (BERTopic [18],
Figure B.1). To capture the language and behavioral patterns typical of the at-risk population, the topic model was trained exclusively on the textual content from the Attempted users’ channels, comprising 757,849 text segments, yielding 166 topics (see Definition 2.2). This model was then used to infer topic distributions across the videos of all research groups (Figure 3). These topic distributions are used in bottom-up and hybrid approaches.
Next, we defined the reference event of each group: a suicide attempt for Attempted (During), a major life event for Control (Major Life Event), and a synthetic event (the midpoint of uploads) for Attempted (Before) and Control (Matches). Topic probabilities were then averaged separately for the pre- and post-event periods. Although the attempt narratives were included in the topic model, they were excluded from subsequent statistical analyses to avoid bias, as they do not represent the typical content of the channels.
@@
In the hybrid approach, we asked a clinical expert: What topics are known risk factors of suicide? which reduced the topics from 166 to 19. In the bottom-up approach, we asked: What topics may serve as indicators of the period before a suicide attempt?, and used stepwise logistic regression to the pre-event period, coding Attempted (During) as 1 and controls as 0, across three control populations and their combined pool, yielding four binary models and narrowing topics from 166 to five.
Following topic selection, we conducted a two-way mixed-effects ANOVA [4], on the time-aggregated topic probabilities, in which the group type was the between-subjects factor, and time (pre- or post-reference point) was the within-subjects factor. Based on the significant effects identified in the ANOVA, we performed Tukey’s post-hoc tests [1], to elucidate group differences and temporal patterns. See §C.3 for detailed statistical analysis.
As a complementary analysis, we extended our time-aggregated approach by conducting two types of longitudinal topic analyses to explore temporal patterns of significant topics. The first was a within-group analysis, where we explored changes over time within the Attempted (During) group. This analysis used $t=0$, allowing us to provide a standardized baseline from which to compare temporal changes. The second was a between-group analysis, in which we compared the Attempted (During) group to the three control groups at each timestep.
Finally, motivated by traditional clinical assessment practices, we conducted a top-down analysis in the form of a psychological factor analysis, aiming to assess whether the behavioral signals identified in our study align with established clinical knowledge on suicidality. This analysis is based upon assessments conducted by a clinical psychologist on the Attempted groups (§2.1.1) and involved one-hot encoding of the psychological factors, then performing a stepwise logistic regression analysis [38], where the outcome variable was coded as 1 for Attempted (During) and 0 for Attempted (Before).

### v2

@@ Abstract
Suicide remains a leading cause of death in Western countries. As social media becomes central to daily life, digital footprints offer valuable insight into suicidal behavior. Focusing on individuals who attempted suicide while uploading videos to their channels, we investigate: How do linguistic patterns on YouTube reflect suicidal behavior, and how do these patterns align with or differ from expert knowledge?
We examined linguistic changes around suicide attempts and compared individuals who attempted suicide while actively uploading to their channel with three control groups: those with prior attempts, those experiencing major life events, and matched individuals from the broader cohort. Applying complementary bottom-up, hybrid, and expert-driven approaches, we analyzed a novel longitudinal dataset of 181 suicide-attempt channels and 134 controls.
In the bottom-up analysis, LLM-based topic-modeling identified 166 topics; five were linked to suicide attempts, two also showed attempt-related temporal changes (Mental Health Struggles, $OR=1.74$; YouTube Engagement, $OR=1.67$; p<.01p<.01) [^2] Odds ratios (OR) quantify the change in the odds of topic prevalence after versus before the suicide attempt within the group of YouTubers who attempted suicide..
In the hybrid approach, clinical experts reviewed LLM-derived topics and flagged 19 as suicide-related. However, none showed significant effects beyond those identified bottom-up. YouTube Engagement, a platform-specific indicator, was not flagged, underscoring the value of bottom-up discovery. A top-down psychological assessment of suicide narratives revealed differing motivations: individuals describing prior attempts aimed to help others ($\beta=-1.69$, $p<.01$), whereas those attempted during the uploading period emphasized personal recovery ($\beta=1.08$, $p<.01$).
@@ 1.  Introduction
Suicide remains one of the most significant public health challenges, particularly among young people in western countries, where it consistently ranks as one of the leading causes of death [naghavi2019global, glenn2020annual]. Traditionally, research on suicidality has focused on factors such as mental health issues (e.g., depression), social isolation, and family dynamics [fergusson2000risk, pelkonen2003child, franklin2017risk], as well as short-term affective fluctuations preceding suicidal behavior [bagge2017trajectories]. These factors were predominantly predefined by experts and were typically examined by relying on patients’ self-reports. However, as social media becomes increasingly prevalent and suicide rates continue to grow, established risk factors and their manifestation are changing, requiring further investigation into suicidal behavior in digital spaces [castillo2020suicide, rabani2023detecting].
@@
Social media have become integral to communication and social development, especially among adolescents [o2011impact]. Many young people turn to these platforms to share their experiences and seek support [fu2013responses, lissak2024colorful], potentially providing a digital footprint, documenting the nuances of everyday life, including moments of joy and distress [de2013predicting].
@@
The intersection of suicide, social media, and Artificial Intelligence (AI) has gained traction in recent years, with much research utilizing natural language processing (NLP), and especially Language Models (LM), to analyze large volumes of social media data [ophir2020deep, lissak2024bored]. However, a common limitation is a reliance on unvalidated suicide risk assessments, such as identifying risk based on messages like "I want to kill myself" or user participation in suicide-related sub-reddits, which introduces biases in both data collection and reliability. Additionally, many studies lack model transparency and have inadequate control groups. Perhaps most importantly, they often fail to incorporate longitudinal analyses that are essential for capturing temporal shifts in relation to suicidal behavior.
@@
Although social media platforms (e.g., Facebook or Reddit) are commonly used to study suicidality, video-sharing platforms like YouTube remain underexplored. As one of the most widely used multimedia-sharing platforms, YouTube enables users to communicate visually and verbally, providing a more direct and unmediated expression. Moreover, YouTube plays a significant role in both shaping and reflecting societal trends related to mental health and suicidality [allgaier2020science, khan2022researching]. Still, a search on PubMed revealed only 28 articles containing the words "suicide" and "YouTube" in their title or abstract, with most focusing on the impact on viewers rather than on the creators [dagar2020high, khasawneh2020examining].
Figure 1: An illustration of our research framework, showing the methodological approaches used in this study. Each circle represents a different approach, with the specific methods listed inside. The computational approach includes Language Models (LM) based
@@
topic modeling to identify behavioral patterns on YouTube (i. §3.1), analyzed through a psychological lens in the discussion section (§4). The domain-expert approach involves a clinical assessment of suicide attempt narratives of predefined psychological factors (ii. §3.3). In the hybrid approach, a domain expert reviewed all the LLM-derived topics and flagged those that may be considered as suicide risk factors, based on current clinical knowledge (iii. §3.4).
@@
This study addresses existing research gaps by bridging the understanding of suicidality in offline and online contexts, considering both behavioral indicators and self-reported experiences, and by incorporating longitudinal analysis of the development of behaviors and feelings before and after suicide attempts. By applying three complementary approaches, as described in Figure 1 and leveraging the strengths of both computational and domain-expert frameworks, we aim to explore our primary research question: How do linguistic patterns on YouTube reflect manifestations of suicidal behavior, and how do these patterns align with or differ from expert knowledge?
@@
To allow this investigation, we first constructed a novel dataset of YouTube channels belonging to individuals who had attempted suicide (Figure 2), with cases verified as life-threatening by a clinical psychologist (Figure A.1). We focused on individuals who attempted suicide after establishing their YouTube channel, referred as Attempted (During). Alongside the Attempted (During) group, which we defined as the treatment group, we gathered three control groups representing varying levels of suicide risk.
@@
Figure 2: The research groups are categorized into treatment - Attempted (During) - and three control groups - Attempted (Before), Control (Matches) and Control (Major Life Event). Both Attempted groups refer to those who attempted suicide during or before their channel upload period, while the other groups refer to control-based matches from the general population, or to those who experienced a major life event during their upload period (and didn’t attempt suicide as far as we know). The sample size for each group is indicated below their respective categories. See data collection pipeline in Figure A.1.
@@
In the bottom-up approach, we employed LM-based topic modeling and analyzed the topic distributions of each video to identify behavioral patterns on YouTube. By examining differences in pre-suicide attempt average topic probability in the treatment group compared to the control groups, we identified potential indicators of suicidal behavior. Then we performed a longitudinal analysis on the identified topics, revealing temporal trends before and after suicide attempts.
Through this analysis, we answer the first part of the research question - how suicidal behaviors manifest on YouTube, and partially address the second part, how are they similar or different from the clinical literature, by examining our findings in light of previous suicide research (see Discussion §4).
@@
Our bottom-up, fully automated computational analysis identified five topics as potential pre-attempt indicators, two of which also exhibited significant temporal shifts around the attempt (§3.1): "Mental Health Struggles" and "YouTube Engagement" (see topics’ representations: Figure 4). "Mental Health Struggles" includes discussions about suicide, self-harm, and depression, aligned with established risk factors. “YouTube Engagement”, which captures typical platform behavior such as audience interaction and content promotion, may serve as an indicator of “normal” YouTube activity, reflecting routine channel maintenance [zhang2023exploring, chung2023really, aquino2024youtube]. Notably, this topic appeared at lower levels in the treatment group prior to the suicide attempt—possibly signaling abnormal behavior and, therefore, a potential risk indicator. Both topics exhibited significantly higher prevalence following the suicide attempt, which we discuss at §4.
Additional three topics emerged as pre-attempt indicators and were not associated with a temporal pattern: (1) "LGBTQ+ Identity and Acceptance", (2) "Rape and Sexual Assault", and (3) "Fashion and Style", detailed in C.3 and discussed in §4. The first two align with established suicide literature, reflecting the heightened distress faced by LGBTQ+ individuals due to discrimination and rejection [sutter2016discrimination], and by survivors of sexual violence due to trauma-related mental health challenges [dworkin2022associations]. Both topics were more prevalent in the attempted groups compared to the other controls. In contrast, Fashion and Style, a less familiar factor in suicide research, showed significantly lower prevalence in the attempted groups, potentially reflecting diminished engagement with self-expression or societal norms.
Moreover, the top-down domain-expert approach showed strong alignment in risk and resilience factors between YouTubers who attempted suicide before and during their upload period, reducing suicide-related tendencies confounders in the comparison between these groups (e.g., depression, family support). The only significant difference was the reason for sharing their attempt story: those who attempted suicide before emphasized their Desire to Help Others, while those who attempted during framed it as part of their Personal Recovery Process (Table C.1).
@@ 2.  Methods
This section outlines the methods used in this study, beginning with data collection and followed by a description of the procedures (Figure 3). The data collection2.1.1) involved constructing a textual dataset comprising four types of YouTube channels creators (Figure 2): belonging to creators who attempted suicide (before or during their YouTube activity), those who experienced a major life event, and matched control users from the general population. We searched YouTube for relevant channels, transcribed the videos and collected channel information, annotated demographic and psychological assessments, and matched them to control channels using a multi-step approach. The procedures2.2) included applying an LM-based topic modeling to the channels’ transcripts and implementing statistical procedures to address our research objective: understanding how suicidal behavior manifests on YouTube.
@@ 2.1.1 Data collection
YouTube served as our primary data source. We identified users who uploaded a video sharing either their own suicide attempt story or a major life event using the YouTube API [^3] API Documentation: developers.google.com/youtube/v3/docs. These stories were typically not the main theme of the channels, but were often shared to raise awareness of suicide or loss. Additionally, we included a control group that matched the treatment users to facilitate a fair comparison with the general population. The data collection procedure is outlined in Figure A.1 and described in more detail below, with full documentation provided in Appendix (A.1 Search, A.2 Pre-processing, A.3 Human Evaluation, A.4 Matching).
@@
We then conducted a human evaluation involving two levels of expertise. First, non-expert channel validation to ensure that only personal channels were included, while excluding educational, news, and family channels. Following this, we extracted user and event information from the attempt and major life event stories (e.g., attempt date). Second, a psychological assessment of the attempt narrative videos of YouTubers who reported a suicide attempt for two purposes: (1) validating the intent of the attempt, and (2) comparing Attempted (During) and Attempted (Before) to suicide-related tendencies to estimate potential confounding factors (e.g., mental illness, risk groups) and to perform a suicide-related factor analysis. See Tables A.3, A.4 for inter-annotator agreement which shows overall substantial to excellent reliability over a subset of examples.
Next, we applied a multi-step matching process (Figure A.3 visualizes the matching algorithm), which allowed us to first search for control cases from the general YouTube population and later match them to users who attempted suicide, based on demographic and channel characteristics.
Finally, we employed data processing, which included collecting the users’ channel videos and metadata (e.g., likes, description) and performing automatic transcription using Whisper [radford2023robust], the current speech-to-text state-of-the-art model.
@@ 2.1.2 Study population
We categorized channels into two primary groups: treatment and control (Figure 2). The treatment consisted of YouTubers who attempted suicide during their upload period, while the first control group included YouTubers who attempted suicide before their upload period. These groups were labeled as Attempted (During) and Attempted (Before). The additional control groups comprised YouTubers who experienced a major life event during their upload period, labeled Control (Major Life Event), and YouTubers matched to the treatment group based on demographic and content characteristics, labeled Control (Matches). The inclusion of three control groups was designed to enable both temporal and cross-sectional comparisons. The Attempted (Before) group allows isolation of trait-level vulnerability effects of suicide, the Major Life Event group controls for non-suicidal distress, and the Matched control group captures baseline platform behavior. Together, these groups facilitate the disentanglement of suicide-specific, temporal, and contextual patterns. See Table A.1 for inclusion and exclusion criteria.
The Final Sample (Table A.2): Of the initial 303 YouTubers in the Attempted group, 104 users were excluded due to non-suicidal cases (e.g., clickbait-like videos like "Bust Myths About Suicide"), non-English content, or an insufficient number of videos, and 18 users were reclassified as Control (Major Life Event) due to the suicide attempt of a close relative. The final set of attempted channels consists of 181 channels (66% female, average age = 25) who uploaded 34,407 videos ($M=190$, $Mdn=83$, $SD=286$). This dataset was later divided into two groups: 42 Attempted (During), defined as the treatment, and 139 Attempted (Before), defined as one of the three controls, based on whether the attempt occurred during or before their upload period (Figure 2).
@@
The final Control (Matches) group included 95 channels (81% female, average age = 23) that met the inclusion criteria and were found as most similar to the 42 Attempted (During) group according to the matching algorithm (Figure A.3). These 95 users collectively uploaded 14,860 videos ($M=182$, $Mdn=95$, $SD=223$).
@@ 2.2 Procedures
Figure 3: The methodological pipeline consists of three approaches: 1. Computational Bottom-Up - Factors were generated using an LM-based topic modeling algorithm, and topic relevance was evaluated using a bottom-up approach; 2. Domain-Expert Top-Down - Factors were predefined by a clinical expert and assessed solely based on the suicide attempt narratives, and a 3. Hybrid Approach - A set of 166 LM-based topics was reduced to 19 based on their relevance to suicide as determined by a clinical expert. The bottom-up and hybrid approaches were also followed by a two-way mixed effect ANOVA examining both temporal and group effects, while the top-down approach was limited to group differences.
Topic Definition A topic is a cluster of semantically related text segments gathered by an unsupervised clustering method of sentence embeddings. Each topic is automatically named by an LLM considering top-associated words and representative texts.
First, we applied an LM-based topic modeling technique using BERTopic [grootendorst2022bertopic], which combines transformer-based sentence embeddings with density-based clustering. The pipeline uses Sentence-BERT [reimers2019sentence] to encode text segments into semantic representations, followed by HDBSCAN clustering to identify thematic groups. Each cluster is then automatically labeled by Llama-2-7B [touvron2023llama] based on representative segments and keywords. We selected this approach because it enables stable, interpretable, and context-aware topic discovery across very large and unstructured datasets with no any earlier assumptions on the topics. It outperforms traditional topic models (e.g., LDA; blei2003latent) and other LLM-based topic modeling because it captures contextual semantic relationships while maintaining computational efficiency and topic stability at scale. Unlike traditional word co-occurrence methods, it detects semantically nuanced themes; unlike other LLM-based approaches, it systematically covers the full corpus and remains more computationally efficient and reliable for comprehensive topic extraction at scale without over-generalizing dominant themes [yang2025enhancing, mu2024large, zhang2022neural]. This method aligns with our goal of identifying non-explicit manifestations of suicidality in over 750,000 YouTube transcripts, where interpretability, stability, and comprehensive coverage are essential (Figure B.1 provides an overview).
To capture the linguistic patterns typical of the at-risk population, the topic model was fitted exclusively on the textual content from the Attempted users’ channels, comprising 757,849 text segments, yielding 166 topics (see Definition 2.2). This model was then used to infer topic distributions across the videos of all research groups (Figure 2). These topic distributions are used in bottom-up and hybrid approaches. Full technical details are provided in Appendix B.1 (prompts, models’ hyper-parameters, and quality metrics over topics).
Next, we defined the reference event of each group: a suicide attempt for Attempted (During), a major life event for Control (Major Life Event), and a synthetic event (midpoint of uploads) for Attempted (Before) and Control (Matches). Topic probabilities were averaged separately for the pre- and post-event periods. Although the attempt narratives were included in the topic model, they were excluded from subsequent statistical analyses to avoid bias, as they do not represent the typical content of the channels.
@@
In the hybrid approach, we asked a clinical expert: What topics are known risk factors of suicide? which reduced the topics from 166 to 19. In the bottom-up approach, we asked: What topics may serve as indicators of the period before a suicide attempt?, and used stepwise logistic regression to the pre-event period [pace2009stepwise], coding Attempted (During) as 1 and controls as 0, across three control populations and their combined pool, yielding four binary models and narrowing topics from 166 to five.  [^4] All stepwise regression findings were replicated using LASSO-regularized logistic regression, confirming the robustness of the identified predictors.
Following topic selection, we fit generalized linear mixed models (GLMMs) [bagiella2000mixed] with a Beta likelihood and a logit link function [ferrari2004beta] to model topic probabilities. In total, we ran $24$ models, one per topic (hybrid: 19 topics, bottom-up: five). For each topic, we calculated the mean probability for the pre-event and post-event periods. The models included fixed effects for Time (After vs. Before the event), Group (see Figure 2), their interaction (Group × Time), and general control variables (Gender, Age, Ethnicity, and Upload activity). The Attempted (During) group and the pre-event period (Before) served as reference categories. We applied Benjamini-Hochberg (BH) FDR correction [benjamini1995controlling] across topics. The complete model specification, including the mathematical formulation and interpretation of coefficients, is provided in Appendix D.1.
As a complementary analysis, we extended our time-aggregated approach by conducting two types of longitudinal topic analyses to explore temporal patterns of significant topics. The first was a within-group analysis, where we explored changes over time within the Attempted (During) group. This analysis used $t=0$, allowing us to provide a standardized baseline from which to compare temporal changes. The second was a between-group analysis, in which we compared the Attempted (During) group to the three control groups at each timestep. Appendix B.2 and B.3 provide further in-depth description of the temporal alignment and testing procedure.
To ensure that our findings were not driven by external or methodological artifacts, we conducted a series of robustness and validation analyses. These analyses evaluated potential confounding factors at multiple levels: (1) a sensitivity analysis excluding ambiguous cases of suicidal ideation to assess the stability of group effects; (2) a control measures analysis examining content shifts during the COVID-19 pandemic; (3) an engagement metrics analysis comparing aggregate audience interaction indicators (e.g., likes) across groups; and (4) an activity analysis assessing individual upload behavior and frequency before and after the reference event. These analysis together with the general control variables included in the GLMM model reduces the possibility of confounding effects in our results (full details in Appendix D.2).
Finally, motivated by traditional clinical assessment practices, we conducted a top-down analysis in the form of a psychological factor analysis, aiming to assess whether the behavioral signals identified in our study align with established clinical knowledge on suicidality. This analysis is based upon assessments conducted by a clinical psychologist on the Attempted groups (§2.1.1) and involved one-hot encoding of the psychological factors, then performing a stepwise logistic regression analysis, where the outcome variable was coded as 1 for Attempted (During) and 0 for Attempted (Before). 4